OpenStudy (anonymous):

Use mental mathematics to find the value of x. 7(x + 5) = 14

OpenStudy (anonymous):

ans: -3 ?

OpenStudy (anonymous):

thanks! i couldnt figure out what i did wrong! i was only off by one number

OpenStudy (anonymous):

7(x + 5) = 14 then, (x+5) = 14/7 then, (x+5)=2 then, x=2-5 finally, x=-3